OpenStudy (anonymous):

Solve the system by substitution. 2x+y=-11 3x-4y=11

jimthompson5910 (jim_thompson5910):

solve the first equation for y 2x+y=-11 y = -11 - 2x ---------------------------------- then plug this into the second equation 3x-4y=11 3x-4(-11 - 2x)=11 3x + 44 + 8x = 11 11x + 44 = 11 I'll let you finish

OpenStudy (anonymous):

would it be (3,5)?

jimthompson5910 (jim_thompson5910):

what do you get when you solve 11x + 44 = 11 for x

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

x = -3 use this to find y

OpenStudy (anonymous):

how do i find y?

jimthompson5910 (jim_thompson5910):

remember we solved for y back in the first equation

jimthompson5910 (jim_thompson5910):

so use y = -11 - 2x and plug in x = -3

OpenStudy (anonymous):

I'm so confused

jimthompson5910 (jim_thompson5910):

y = -11 - 2x y = -11 - 2(-3) ... replace x with -3 (since x = -3) y = ???

OpenStudy (anonymous):

-5?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so, would it be (-3,-5) or (-5,-3)?

jimthompson5910 (jim_thompson5910):

hint: all ordered pair points are of the form (x,y)

OpenStudy (anonymous):

You got -3 for x. so in the x position you put -3... you have y = -11 - 2(-3).... this becomes y = -11+ 6, or y= -5.... you have x = -3 and y = -5, ordered pairs are written in (x,y) form... so you have (-3, -5) as your ordered pair